# Effect of the depolarizing field on the domain structure of an improper ferroelectric

### v2

@@ Abstract
We show that, contrary to common belief, the depolarizing electric field generated by bound charges at thin-film surfaces can have a substantial impact on the domain structure of an improper ferroelectric with topological defects. In hexagonal-manganite thin films, we observe in phase-field simulations that through the action of the depolarizing field, (i) the average magnitude of the polarization decreases, (ii) the local magnitude of the polarization decreases with increasing distance from the domain walls, and (iii) there is a significant alteration of the domain-size distribution, which is visualized with the pair-correlation function. We conclude that, in general, it is not appropriate to ignore the effects of the depolarizing field for thin film ferroelectrics.
@@ I Introduction
In this work, we predict that even in improper ferroelectrics and despite the presence of topological defects pinning the domains the depolarizing field can have an unexpected, significant impact on the domain structure. We perform phase-field simulations of free-standing thin films of hexagonal manganites, a lattice-distortively driven improper ferroelectric, and show that the average magnitude of the spontaneous electric polarization is lowered by the depolarizing field. Remarkably, the magnitude of the polarization drops with increasing distance from the domain walls. Furthermore, the structure factor and the cross-correlation function reveal a significant effect on the domain-size distribution. Both change from a 2D-Gaussian function in the absence of a depolarizing field to donut-like distribution in its presence. This demonstrates that it is not appropriate to ignore the depolarizing field in thin film improper ferroelectrics, regardless of topological defects.
@@ II The Landau expansion of hexagonal manganites
where the electric field $E$ is obtained from the Gauss law, choosing a value of $\varepsilon=0.5$ to make the electrostatic effects more evident. As shown in the supplemental material, increasing the dielectric constant to a realistic value of $\varepsilon=12.5$ [12, 10] decreases the effect of the depolarizing field but does not affect the results qualitatively. The purpose of the choice of $\varepsilon=0.5$ is to enhance the magnitude of these effects, which permits reducing the size of our computational mesh while keeping the required computational time within feasible limits.
The model in Eq. 1 is derived for bulk crystals. Here, we adapt it to the description of thin films by using periodic boundary conditions in the in-plane directions and open boundary conditions in the out-of-plane direction, as previously done in Ref. [9]. Furthermore, we assume that the system is a perfect insulator and that it is devoid of any free charges. We choose open boundary conditions over closed-circuit boundary conditions to to bring out the effects of the electrostatic interaction as clearly as possible and to avoid the significant complexity and computational time of simulating free charge carriers and charge currents [11].
@@
where $\eta$ is an order parameter, here $Q$, $\Phi$, or $\mathcal{P}$. The expression $\delta/\delta\eta$ corresponds to a functional derivative. The electrostatic field of the system is computed from the Gauss law [10]. Finally, we compute the structure factor of the polarization-density field by taking the square of the Fourier transform of the system, in equivalence to scattering experiments. The structure factor then gives us insight on the domain-size distribution. The pair correlation of the system is computed from an inverse transform of the structure factor to derive typical domain sizes. A detailed description of the computational details and data analysis can be found in Appendix A.
@@ III Results
The electrostatic interaction can be divided into two contributions. Bound charges at the surface of the thin film give raise to the depolarizing field, which is distinguished from the electric field generated by bound charges at domain walls. We treat the effect of the depolarizing field in Figs. 2-4. The effect of bound charges at domain walls in treated in Figs. 5 and 6
@@
Figure 2: Dependence of improper-ferroelectric order on depolarizing field, visualized with phase-field simulations in real space. (a) Domain pattern for a simulation that ignores the electrostatic interaction. (b) Domain pattern for a simulation that includes the electrostatic interaction assuming a dielectric constant $\varepsilon=0.5$. (c) Close-up on the green paths through panels (a) (left) and (b) (right). (d) Domain size along the green path for the simulations without (top) and with (bottom) the depolarizing field. (e, f) Polarization profile along a section, shown as yellow line in (a, b). The gradient and resulting dip in polarization is clearly visible in (f).
@@
In Fig. 4, the structure factor and the pair correlation of the polarization of has been computed by summing the result of twenty simulations with different initial fields. This computation is equivalent to the simulation of a scattering experiment, and its technical details are described in Appendix A. Figures 4a, 4b, and 4c show the exponentially decaying structure factor and pair correlation of a system that does not include the electrostatic interaction. The structure factor is indicative of a broad domain-size distribution that is typical for hexagonal manganites in bulk [17] or for thin films where the depolarizing field is fully screened [18, 16]. This contrasts strikingly with the structure factor of simulations that include the electrostatic interaction. The observed ring in Fig. 4d and the pair correlation in Figs. 4e and 4f, which takes the form of a spherical wave, are indicative of a peak in the domain-size distribution, a stark difference from the case without the electrostatic interaction. The value of this peak is given by the diameter of the red disks in Fig. 4. This change in the domain patterns is further illustrated in real space in Fig. 2c, which shows an illustrative path exemplifying the change in domain size, and in and in Fig. 2d, which shows the change in domain-size distribution along this path.
Figure 4: Structure factor and pair correlation of twenty combined simulations for (a-c) systems without electrostatic interaction and (d-f) systems with electrostatic interaction. (a) Structure factor and (b,c) pair correlation with two color scales. (d-f) Same as (a-c), but with electrostatic interaction for $\varepsilon=0.5$. If electrostatic interaction and hence a depolarizing field is present, the ring-like structure factor in (d) and the spherical wave in (e) and (f) indicate a narrow domain size distribution with a clear peak, the center of the distribution given by the diameter of the red disk in (f).
@@ IV Discussion
In Figs. 2 to 5, we have shown that there are three are a number of effects that are caused by the depolarizing field exerted by bound charges at surfaces and interfaces of thin films. (i) The averaged magnitude of the polarization is lowered, (ii) the polarization declines away from the domain walls with a dip in polarization in the approximate center of the domain, and (iii) a flat domain size distribution for simulations where the depolarizing field is ignored contrasts with a distribution with a clear peak for simulations that include the depolarizing field. Furthermore, due to the electric field generated by bound charges at domain walls, domain walls have a stronger tendency to align parallel to the z-axis if the electrostatic energy is considered, as shown in Figs. 5 and 6.
Note that modification of domain-size is even more surprising considering that the ferroelectricity in hexagonal manganites is not only improper, but also exhibits topological defects. Despite this two-fold opposition to the action of the depolarizing field, it manifests itself with unprecedented clarity in the structure factor and cross-correlation. Our work shows that the depolarizing field can have a significant effect on the domain structure of improper ferroelectrics, even if topological defects are present, and that, in general, it is not appropriate to ignore it.
@@ V Conclusion
In this work, we have shown that the depolarizing field has a significant effect on the domain pattern of both the primary lattice-distortive and the secondary ferroelectric order in thin film hexagonal manganites. We find that the average magnitude of the polarization in the domains is reduced, the system exhibits local variations in polarization, and there is a significant change in the domain-size distribution. Furthermore, the energy penalty of charged domain walls causes the domain walls to align parallel to the direction of the polarization.
The emergence of these effects shows that although the primary order and the topological defects are still the main driving force behind the domain structure, the effect of the depolarizing field is significant and non-negligible. We have also shown these effects are inconspicuous and only reveal themselves clearly in reciprocal space, in contrast to the domain structure of the topological defects which is apparent in real space. This may explain why these effects have not been considered in previous work.
@@ Acknowledgements.
We acknowledge useful discussions with Nicola Spaldin, Morgan Trassin, Arkadiy Simonov, and Quintin Meier. This work was funded by the Swiss National Science Foundation (SNSF) through grant numbers 200021_178825 and 200021_215423.
@@ Appendix A Detailed Numerical Methods
is added, where $P=\mathcal{P}\cdot 9.031\,e/365.15\,\text{\AA}^{3}$ is the polarization of the system [5] and $e$ is the elemental charge. The electric field $E$ is computed via the Gauss law. We assume that our system is a perfect insulator, and hence the total charge density is given by bound charges only according to $\rho=-\nabla P$. From the charge density, the electrostatic potential is computed via the Poisson equation $\Delta\varphi=-\rho/\varepsilon$. Finally, from the electrostatic potential, the electric field can be obtained via $E=-\nabla\varphi$.
@@
We initialize the order parameters $Q_{x}=Q\,\text{cos}\Phi$, $Q_{y}=Q\,\text{sin}\Phi$, and $\mathcal{P}$ with a random, uniform distribution on the interval [-0.1, 0.1]. Qualitatively, the domain pattern is independent of the initial random configuration. Quantitatively, however, the specific domain pattern depends on the initial configuration. The Ginzburg-Landau equations are then integrated in a finite difference scheme with a Runge-Kutta 4 integrator. For the main results in Figs. 2 to 4, we simulate a system of size $n_{x}\times n_{y}\times n_{z}=128\times 128\times 3$ with lattice spacing $\Delta x=\Delta y=0.1\,\text{nm}$ and $\Delta z=0.3\,\text{nm}$. Although phase-field simulations are a freely scalable model, it is still sensible to choose a physically meaningful scale for the simulations. Here, the thickness of the simulated film in $z$-direction has been chosen to approximately correspond to a single monolayer, which is the limit of the physically meaningful range. With this choice, all domain walls are parallel to the z-direction, and there are no charged domain walls. Bound charges only accumulate at the surfaces of the thin films and all electrostatic effects stem from the depolarizing field. We use a time step of $\Delta t=5\cdot 10^{-4}$. The dielectric constant of the material $\varepsilon$ is chosen as described in the main text and the supplemental material. The system was iterated for $n=10^{4}$ time steps. The Poisson equation to obtain the electrostatic potential is solved using a custom V-multigrid solver using Jacobi iterations. Simulations of thick films in Fig. 5 have been performed with system of size $n_{x}\times n_{y}\times n_{z}=128\times 128\times 15$ with lattice spacing $\Delta x=\Delta y=0.2\,\text{nm}$ and $\Delta z=0.3\,\text{nm}$ and with a time step of $\Delta t=5\cdot 10^{-4}$. Simulations including the electrostatic energy were iterated for $n=10^{5}$ time steps, and a dielectric constant of $\varepsilon=0.5$ was used. The system that ignores electrostatic interactions was iterated for $n=4.3\cdot 10^{4}$ time steps to obtain domains of similar size. For the quantitative analysis of thick films in Fig. 6, simulations with system of size $n_{x}\times n_{y}=128\times 128$ with variable $n_{z}$ have been performed, with lattice spacing $\Delta x=\Delta y=0.2\,\text{nm}$ and $\Delta z=0.3\,\text{nm}$. For the thinnest films, we chose $n_{z}=3$, which approximately corresponds to the thickness of a single monolayer. The time step has been set to $\Delta t=5\cdot 10^{-4}$. The system was iterated for $n=10^{5}$ time steps for simulations that include and that ignore the electrostatic interaction. For the simulations with electrostatic interaction, we have set the dielectric constant to $\varepsilon=0.5$.
Data analysis. We compute the structure factor of the system by first taking a slice of the computational mesh of the thin film at $z=0$. Then, we compute the two-dimensional Fourier transform of the domain pattern and square the absolute values of the Fourier components. Fourier transformations of twenty simulations are combined to form the structure factor of the system. To compute the pair correlation, we first set the Fourier component of the structure factor $S$ at zero wave vector to zero to remove a constant offset value, that is, $S(\mathbf{k}=0)=0$. The pair correlation is obtained by then transforming the structure factor back to real space.

### v3

@@ Abstract
We show that, contrary to common belief, the depolarizing electric field generated by bound charges at thin-film surfaces can have a substantial impact on the domain structure of an improper ferroelectric with topological defects. In hexagonal-manganite thin films, we observe in phase-field simulations that through the action of the depolarizing field, (i) the average magnitude of the polarization decreases, (ii) the local magnitude of the polarization decreases with increasing distance from the domain walls, and (iii) there is a significant alteration of the domain-size distribution and average domain size, which is visualized with the pair-correlation function. We conclude that, in general, it is not appropriate to ignore the effects of the depolarizing field for thin film ferroelectrics.
@@ I Introduction
In this work, we predict that even in improper ferroelectrics and despite the presence of topological defects pinning the domains the depolarizing field can have an unexpected, significant impact on the domain structure. We perform phase-field simulations of free-standing thin films of hexagonal manganites, a lattice-distortively driven improper ferroelectric, and show that the average magnitude of the spontaneous electric polarization is lowered by the depolarizing field. Remarkably, the magnitude of the polarization drops with increasing distance from the domain walls. Furthermore, the Fourier-transformed intensities and the cross-correlation function reveal a significant effect on the domain-size distribution. Both change from a 2D-Gaussian function in the absence of a depolarizing field to donut-like distribution in its presence. This demonstrates that it is not appropriate to ignore the depolarizing field in thin film improper ferroelectrics, regardless of topological defects.
@@ II The Landau expansion of hexagonal manganites
where the electric field $E$ is obtained from the Gauss law
$\nabla E=-\frac{\nabla P_{\mathrm{s}}}{\varepsilon_{\rm b}\varepsilon_{0}}.$ (3)
Here, $\varepsilon_{0}$ is the vacuum permittivity and $\varepsilon_{\rm b}$ is the background dielectric constant [12, 13]. The background dielectric constant contains the electronic dielectric response and the dielectric response of all normal modes in the system with the exception of the mode describing the polarization $P$. Here, we choose a background dielectric constant of $\varepsilon_{\rm b}=1$ to make the electrostatic effects more evident. As shown in the supplemental material, increasing the background dielectric constant to the experimental value of $\varepsilon_{\rm b}=8.9$ [14, 10] merely decreases the effect of the depolarizing field but does not affect the results qualitatively. On the practical side, choosing the value $\varepsilon_{\rm b}=1$ permits us to reduce the size of our computational mesh while keeping the required computational time within feasible limits.
The model in Eq. 1 is derived for bulk crystals. Here, we adapt it to the description of thin films by using periodic boundary conditions in the in-plane directions and open boundary conditions in the out-of-plane direction, as previously done in Ref. [9]. Furthermore, we assume that the system is a perfect insulator and that it is devoid of any free charges. We choose open boundary conditions over closed-circuit boundary conditions to bring out the effects of the electrostatic interaction as clearly as possible and to avoid the significant complexity and computational time of simulating free charge carriers and charge currents [11].
@@
where $\eta$ is an order parameter, here $Q$, $\Phi$, or $\mathcal{P}$. The expression $\delta/\delta\eta$ corresponds to a functional derivative. The electrostatic field of the system is computed from the Gauss law [10]. Finally, we compute the Fourier-transformed intensity of the polarization-density field by taking the square of the Fourier transform of the system, in equivalence to scattering experiments. The Fourier-transformed intensity then gives us insight on the domain-size distribution. The pair correlation of the system is computed from an inverse transform of the intensity to derive typical domain sizes. A detailed description of the computational details and data analysis can be found in Appendix A. As shown in the supplemental material, the full-width half-maximum of the central peak of the pair correlation is proportional to the average domain size [19].
@@ III Results
The electrostatic interaction can be divided into two contributions. Bound charges at the surface of the thin film give raise to the depolarizing field, which is distinguished from the electric field generated by bound charges at domain walls. We treat the effect of the depolarizing field in Figs. 2-4. The effect of bound charges at domain walls is treated in Figs. 5 and 6.
@@
Figure 2: Dependence of improper-ferroelectric order on depolarizing field, visualized with phase-field simulations in real space. (a) Domain pattern for a simulation that ignores the electrostatic interaction. (b) Domain pattern for a simulation that includes the electrostatic interaction assuming a background dielectric constant $\varepsilon_{\rm b}=1$ (see text on this choice). (c) Close-up on the green paths through panels (a) (left) and (b) (right). (d) Domain size along the green path for the simulations without (top) and with (bottom) the depolarizing field. (e, f) Polarization profile along a section, shown as yellow line in (a, b). The gradient and resulting dip in polarization is clearly visible in (f).
@@
In Fig. 4, the Fourier-transformed intensities and the pair correlation of the polarization has been computed by summing the result of twenty simulations with different initial fields. We further perform data augmentation. This computation is equivalent to the simulation of a scattering experiment, and its technical details are described in Appendix A. Figures 4a, 4b, and 4c show the exponentially decaying intensity and pair correlation of a system that does not include the electrostatic interaction. The intensity is indicative of a broad domain-size distribution that is typical for hexagonal manganites in bulk [20] or for thin films where the depolarizing field is fully screened [21, 18]. This contrasts strikingly with the intensity distribution in simulations that include the electrostatic interaction. The observed ring in Fig. 4d and the pair correlation in Figs. 4e and 4f, which takes the form of a spherical wave, reveal a narrowing of the domain-size distribution, whereas a smaller width of the central peak corresponds to a decrease in the average domain size, an aspect that is further discussed in the supplemental material. This change in the domain patterns is further illustrated in real space in Fig. 2c, which shows an illustrative path exemplifying the change in the shape of the domains, and in and in Fig. 2d, which shows the associated change in domain-size distribution.
Figure 4: Fourier-transformed intensity and pair correlation of twenty combined simulations for (a-c) systems without electrostatic interaction and (d-f) systems with electrostatic interaction. (a) Fourier-transformed intensity and (b,c) pair correlation with two scaling ranges. (d-f) Same as (a-c), but with electrostatic interaction for $\varepsilon_{\rm b}=1$. Note that data augmentation was used for this analysis (see Appendix A). The anisotropy of the plot results from the quadratic simulation mesh and the artifacts its size limitation introduces. If electrostatic interaction and hence a depolarizing field is present, the ring-like intensity in (d) and the spherical wave in (e) and (f) indicate a narrowing of the domain-size distribution, whereas a smaller width of the central peak corresponds to a decrease in the average domain size.
@@ IV Discussion
In Figs. 2 to 5, we have shown that there are three effects that are caused by the depolarizing field exerted by bound charges at surfaces and interfaces of thin films. (i) The averaged magnitude of the polarization is lowered, (ii) the polarization declines away from the domain walls with a dip in polarization in the approximate center of the domain, and (iii) a flat domain size distribution and a large average domain size for simulations where the depolarizing field is ignored contrasts with a distribution with a clear peak and a smaller average domain size for simulations that include the depolarizing field. Furthermore, due to the electric field generated by bound charges at domain walls, domain walls have a stronger tendency to align parallel to the z-axis if the electrostatic energy is considered, as shown in Figs. 5 and 6.
Note that modification of the domain-size shape and distribution is even more surprising considering that the ferroelectricity in hexagonal manganites is not only improper, but also exhibits topological defects. Despite this two-fold opposition to the action of the depolarizing field, it manifests itself with unprecedented clarity in the Fourier-transformed intensity and cross-correlation. Our work shows that the depolarizing field can have a significant effect on the domain structure of improper ferroelectrics, even if topological defects are present, and that, in general, it is not appropriate to ignore it.
@@
As mentioned in Section II, our choice of $\varepsilon_{\rm b}=1$ enhances the effect of the depolarizing field in our simulations without changing these effects qualitatively. In the supplemental material, we show a series of simulations between the physical limit of $\varepsilon_{\rm b}=1$ and the experimental value of $\varepsilon_{\rm b}=8.9$, and it is apparent how the influence of the depolarizing-field decreases with $\varepsilon_{\rm b}$. While it may be argued that the depolarizing-field correction may eventually become negligible, we point out that even a small correction is likely to have an effect on domain formation in the first place, as it occurs at the point of instability when crossing the Curie temperature.
@@ V Conclusion
In this work, we have shown that the depolarizing field has a significant effect on the domain pattern of both the primary lattice-distortive and the secondary ferroelectric order in thin film hexagonal manganites. We find that the average magnitude of the polarization in the domains is reduced, the system exhibits local variations in polarization, and there is a significant change in the domain-size distribution and the average domain size. Furthermore, the energy penalty of charged domain walls causes the domain walls to align parallel to the direction of the polarization.
The emergence of these effects shows that although the primary order and the topological defects are still the main driving force behind the domain structure, the effect of the depolarizing field is significant and, in general, non-negligible. We have also shown these effects are inconspicuous and only reveal themselves clearly in reciprocal space, in contrast to the domain structure of the topological defects which is apparent in real space. This may explain why these effects have not been considered in previous work.
@@ Acknowledgements.
We acknowledge useful discussions with Nicola Spaldin, Morgan Trassin, and Quintin Meier. We further thank Martin Lilienblum for providing the piezoresponse-force-microscopy images shown in the supplementary material. This work was funded by the Swiss National Science Foundation (SNSF) through grant numbers 200021_178825 and 200021_215423.
@@ Appendix A Detailed Numerical Methods
is added, where $P=\mathcal{P}\cdot 9.031\,e/365.15\,\text{\AA}^{3}$ is the polarization of the system [5] and $e$ is the elemental charge. The electric field $E$ is computed via the Gauss law. We assume that our system is a perfect insulator, and hence the total charge density $\rho$ is given by bound charges only according to $\rho=-\nabla P$. From the charge density, the electrostatic potential is computed via the Poisson equation $\Delta\varphi=-\rho/\varepsilon_{\rm b}\varepsilon_{0}$. The background dielectric constant contains all dielectric contributions that are not caused by the spontaneous polarization $P_{\mathrm{s}}$, that is, electronic responses and responses from all normal modes except $\Gamma^{-}_{2}$ [10]. Finally, from the electrostatic potential, the electric field can be obtained via $E=-\nabla\varphi$.
@@
We initialize the order parameters $Q_{x}=Q\,\text{cos}\Phi$, $Q_{y}=Q\,\text{sin}\Phi$, and $\mathcal{P}$ with a random, uniform distribution on the interval [-0.1, 0.1]. Qualitatively, the domain pattern is independent of the initial random configuration. Quantitatively, however, the specific domain pattern depends on the initial configuration. The Ginzburg-Landau equations are then integrated in a finite difference scheme with a Runge-Kutta 4 integrator. For the main results in Figs. 2 to 4, we simulate a system of size $n_{x}\times n_{y}\times n_{z}=128\times 128\times 3$ with lattice spacing $\Delta x=\Delta y=0.1\,\text{nm}$ and $\Delta z=0.3\,\text{nm}$. Although phase-field simulations are a freely scalable model, it is still sensible to choose a physically meaningful scale for the simulations. Here, the thickness of the simulated film in $z$-direction has been chosen to approximately correspond to a single monolayer, which is the limit of the physically meaningful range. With this choice, all domain walls are parallel to the z-direction, and there are no charged domain walls. Bound charges only accumulate at the surfaces of the thin films and all electrostatic effects stem from the depolarizing field. We use a time step of $\Delta t=5\cdot 10^{-4}$. The background dielectric constant of the material $\varepsilon_{\rm b}=1$ is chosen as described in the main text and the supplemental material. The system was iterated for $n=10^{4}$ time steps. The Poisson equation to obtain the electrostatic potential is solved using a custom V-multigrid solver using Jacobi iterations. Simulations of thick films in Fig. 5 have been performed with system of size $n_{x}\times n_{y}\times n_{z}=128\times 128\times 15$ with lattice spacing $\Delta x=\Delta y=0.2\,\text{nm}$ and $\Delta z=0.3\,\text{nm}$ and with a time step of $\Delta t=5\cdot 10^{-4}$. Simulations including the electrostatic energy were iterated for $n=10^{5}$ time steps, and $\varepsilon_{\rm b}=1$. The system that ignores electrostatic interactions was iterated for $n=4.7\cdot 10^{4}$ time steps to obtain domains of similar size. For the quantitative analysis of thick films in Fig. 6, simulations with system of size $n_{x}\times n_{y}=128\times 128$ with variable $n_{z}$ have been performed, with lattice spacing $\Delta x=\Delta y=0.2\,\text{nm}$ and $\Delta z=0.3\,\text{nm}$. For the thinnest films, we chose $n_{z}=3$, which approximately corresponds to the thickness of a single monolayer. The time step has been set to $\Delta t=5\cdot 10^{-4}$. The system was iterated for $n=10^{5}$ time steps for simulations that include and that ignore the electrostatic interaction. For the simulations with electrostatic interaction, we have set $\varepsilon_{\rm b}=1$.
Data analysis. We compute the Fourier-transformed intensity of the system by first taking a slice of the computational mesh of the thin film at $z=0$. We first subtract the average of the polarization from the system and then compute the two-dimensional Fourier transform of the domain pattern.
$F(\mathbf{k})=\mathcal{F}[P(\mathbf{r})-P_{\text{av}}]$ (9)
Here, $F(\mathbf{k})$ corresponds to the structure factor, $\mathcal{F}$ denotes the Fourier transform and $P_{\text{av}}$ denotes the average polarization. We square the absolute values of the structure factor to obtain the Fourier-transformed intensity of the system.
$I(\mathbf{k})=|F(\mathbf{k})|^{2}$ (10)
Fourier-transformed intensities of twenty simulations are combined to obtain better statistics. Furthermore, we perform data augmentation by applying a 90° rotation, a mirror operation, and a combined 90° rotation and mirror operation. The pair correlation is obtained by then transforming the Fourier-transformed intensity back to real space.
$A(\mathbf{r})=\mathcal{F}^{-1}\left[I(\mathbf{k})\right]=\mathcal{F}^{-1}\left[|F(\mathbf{k})|^{2}\right]$ (11)
As shown in the supplemental material, the full-width half-maximum of the pair correlation is proportional to the average domain size of the system.